# GLL: A Differentiable Graph Learning Layer for Neural Networks

### v1

@@ 3.2.1 Sparse and efficient gradient computations
4 Numerical Experiments
In the following section, we showcase the strong performance of GLL in a variety of tasks. In 4.1, we show the geometric effect of varying $\tau$ in Equation (15) in a small neural network on a toy dataset. 4.2 and 4.3 use the standard Laplace learning algorithm (Equation (11)) with a variety of neural network backbones to show the improved generalization and adversarial robustness of GLL. While our theoretical results are for any graph Laplacian-based learning algorithm, we leave numerical experiments on methods like Poisson learning and $p$-Laplace learning for future work.
4.1 Toy Datasets
In this subsection, highlight the differences between the learned embeddings of neural networks when trained on GLL compared to a softmax-based classifier. When trained with GLL, the neural network’s latent space recovers the underlying geometry of the data, which may be desireable for applications. Moreover, we visualize the effect on the decay parameter $\tau$ from Section 2.4. Larger values of $\tau$ result in tighter clustering of the data in the latent space, at the expense of preserving the geometry. The tradeoff between these two will depend on the application; for now, we give a brief example with a toy dataset (two moons (Pedregosa et al. (2011))).
The architecture backbone for this experiment is a multi-layer perceptron neural network with three linear layers. The hidden size is 64 and the output size is two for two dimensional visualization purposes. We choose an initialization where the samples are sufficiently mixed, making the problem more difficult (Figure 2). We arbitrarily choose the first ten samples as the labeled samples from which GLL will propagate the labels. We use cross entropy loss. To make the results more interpretable, we fix the $\varepsilon_{k}$ term to be one. The embeddings for this ablation study is shown in Figure 3. We compare five values of $\tau$: 0, 0.001, 0.01, 0.1, and 0.5. The figure shows the learned embedding of the dataset from the neural network across various training epochs.
We note that this regime is deliberately constrained to two dimensions, meaning the data will struggle to move past each other in the latent space. The seed for these samples was chosen to be difficult and has the a very mixed initialization for the sample embeddings. Despite this, after sufficient training, the model successfully separates the data.
@@ 4.1 Toy Datasets
Figure 2: The initial embedding of the two moons dataset from the random seed. The blue and red correspond to the different classes, with the starred samples corresponding to the labeled samples for the respective classes.
@@
Figure 3: Visualization of effect of $\tau$ on spatial embeddings. Each row corresponds to a graph learning model with a specified value of $\tau$ or a softmax classification head. The columns correspond to the epoch of the training. With low values of $\tau$, the model separates the data by class but resembles the original two moons shape. Increasing $\tau$ forces tighter clustering and localization. The softmax layer simply learns a linear separation with no regard to relational information. Further training with the softmax layer increases the scale, which is not seen without the axes.
These experiments highlight the geometry of the learned embeddings under GLL. We compare these embeddings to the ones learned by the same three layer MLP encoder network but with a projection head and softmax layer in place of the GLL layer. The baseline MLP encoder with the projection head will simply linearly separate the data. The encoder immediately followed by the GLL layer, however, will quickly recover the original two moons shape. The addition of the $\tau$ decay parameter will encourage tighter clustering of the classes around the base labels, rather than a two moons shape. For low values of $\tau$, we still see initial convergence to the two moons, since the driving loss term comes from label predictions. Once the two moons structure is reached, the loss from the prediction is negligible and the regularization term begins to drive the loss, which results in tighter clusters. If the $\tau$ value is too large, then the two moons shape is skipped entirely in favor of a tighter clustering. A large $\tau$ value could also cause the labeled groups to splinter into multiple clusters localized about different base samples.
4.2 Comparison to the Multilayer Perceptron (MLP)
One direct application of our proposed GLL is to serve as a classifier, replacing a projection layer and softmax. In Figure 1, we illustrate the differences between the neural network structure incorporating our GLL and the standard structure. Since GLL does not include any trainable parameters, the reduced number of parameters and layers makes our network structure smoother to train. In this subsection, we design experiments to compare the performance of using an MLP and our GLL as classifiers under the same feature encoder. We conclude that by replacing the MLP with GLL as the classifier in classic network structures, the training speed of the network will be accelerated, and it can effectively avoid problems such as training stagnation caused by vanishing gradients.
We consider two datasets, the FashionMNIST dataset (Xiao et al. (2017)) and the CIFAR-10 dataset (Krizhevsky et al. (2009)), for the experiments in this section. For the FashionMNIST dataset we use a simple CNN-based network (Table 1) with a feature encoder containing 3 convolutional layers (Conv) and 1 fully connected layer (FC). Each conv and FC layer is followed by a corresponding ReLU activation function, which is omitted in the table. On the CIFAR-10 dataset, we employ standard ResNet-18 and ResNet-50 architectures, where the feature encoder consists of several ResNet blocks followed by an FC layer (Table 2). For both datasets, the MLP Classifiers consist of 2 FC layers followed by a softmax layer. To compare with our GLL layer, we keep the feature encoder structure unchanged from Tables 1,2 and replace the MLP classifier with the GLL classifier. The outputs are used to calculate the cross-entropy loss with the ground-truth labels and perform backpropagation and optimization.
We also compare with the neural networks trained with weighted nonlocal Laplacian (WNLL, Wang and Osher (2021)), called WNLL-DNN, on the CIFAR-10 dataset. WNLL-DNN uses a graph-based classifier, but approximates the gradients using a linear layer. We used the accuracy values of ResNet-18 and ResNet-50 reported in their paper. The detailed network architecture for WNLL-DNN can be obtained from their GitHub repository  [^5] https://github.com/BaoWangMath/DNN-DataDependentActivation/tree/master. The only difference between our ResNet architecture (Table 2) and theirs is the number of parameters in the fully connected layers after the ResNet blocks.
@@ 4.2 Comparison to the Multilayer Perceptron (MLP)
Table 1: The network structure used to compare MLP and GLL classifiers for the FashionMNIST dataset(Section 4.2).
Structure Layer Details Feature Size
Input Image - From CIFAR-10 Dataset 32 $\times$ 32 $\times$ 3
Feature Encoder ResNet-18 8 ResNet Blocks $512$
or ResNet-50 $16$ ResNet Blocks $2048$
FC 512/2048 $\rightarrow$ 128 128
MLP Classifier FC1 128 $\rightarrow$ 32 32
FC2 32 $\rightarrow$ 10 10
Softmax dim: 1 10
Table 2: The structure used to compare MLP and GLL classifiers for the CIFAR-10 dataset (Section 4.2).
All experiments in this section are fully supervised with the entire training set of the FashionMNIST and CIFAR-10 datasets. We report results on the following experiments:
  • (FashionMNIST: 100 MLP) Directly training a neural network with an MLP Classifier from scratch. In this case, we simultaneously record the training loss and test accuracy of both the MLP and GLL during the training process. Specifically, the output of the Feature Encoder is passed to both the MLP Classifier and a GLL. Since GLL has no parameters, this does not affect the training. Then, based on the output of this GLL, we can calculate the GLL loss/accuracy during the MLP training process. Here, we train for 100 epochs.
  • (FashionMNIST: 50 GLL) Training a neural network with GLL as the classifier from scratch, only training for 50 epochs.
  • (FashionMNIST: 50 MLP + 50 GLL) Starting from the 50th epoch of the MLP training process, we switch to using the backpropagation gradient information from the GLL Classifier for training and continue training for another 50 epochs.
@@
  • (FashionMNIST: 75 MLP + 25 GLL) Starting from the 75th epoch of the MLP training process, we switch to using GLL for training and continue training for another 25 epochs.
  • (CIFAR-10 MLP and GLL) We pretrain the feature encoder on the CIFAR-10 training dataset using the SimCLR method, which does not require any label information. Subsequently, we use these SimCLR weights as the initial weights for the feature encoder, combined with either an MLP classifier (PyTorch’s default weight initialization) or our GLL classifier (no trainable parameters) to perform fully supervised training. More details about training on CIFAR-10 are in Appendix A.
  • (CIFAR-10 WNLL-DNN) We use the accuracy values reported in the paper (Wang and Osher (2021)).
Dataset Trial Network Epochs Accuracy
FashionMNIST MLP Only Customized 100 MLP 54.58% MLP
85.34% GLL
GLL Only 50 GLL 91.10% GLL
GLL-50 50 MLP + 50 GLL 91.32% GLL
GLL-75 75 MLP + 25 GLL 91.09% GLL
CIFAR-10 MLP ResNet-18 800 MLP 95.42% MLP
GLL 800 GLL 96.10% GLL
WNLL-DNN N/A 95.35% WNLL
MLP ResNet-50 800 GLL 95.77% MLP
GLL 800 GLL 96.54% GLL
WNLL-DNN N/A 95.83% WNLL
Table 3: Test Accuracy comparisons. We present accuracy results on FashionMNIST and CIFAR-10 datasets with fully supervised networks. In Epoch and Accuracy rows, MLP, GLL, or WNLL are the classifiers used to train and evaluate test accuracy. For FashionMNIST, we use our customized network (Table 1), which is intentionally designed with a large number of parameters to make it challenging to train, emphasizing the importance of effective training strategies. We consider four different training strategies: MLP Only, GLL Only, and two MLP-GLL mixed ones, GLL-50 and GLL-75. For CIFAR-10, we consider ResNet-18,50 as the main network structure. where each is followed by an MLP or a GLL classifier for the training. As for the WNLL-DNN, we use the accuracy values reported in the paper (Wang and Osher (2021)). Our GLL Classifier achieves higher accuracy than the MLP classifier in direct or WNLL training.
@@
Figure 4: Training loss and test accuracy comparison of four different training methods on the FashionMNIST dataset, using a fixed feature encoder and either an MLP or GLL as the classifier. The methods include (1) MLP trained from scratch, (2) GLL trained from scratch (GLL-0), (3) GLL trained after 50 epochs of MLP training (GLL-50), and (4) GLL trained after 75 epochs of MLP training (GLL-75). For the MLP classifier, both MLP and GLL training losses and test accuracies are recorded during training, while for the GLL classifier, only the corresponding GLL classifier’s training loss and test accuracy are reported.
Accuracy comparison is shown in Table 3. For the FashionMNIST results, we provide Figure 4, which compares the four training strategies. Figure 4(a) presents the training loss curves as the epochs progress, while Figure 4(b) displays the test accuracy curves over the epochs. Since GLL-50 and GLL-75 are trained after 50 and 75 epochs of MLP training, respectively, their corresponding curves do not start from 0.
From Table 3, it is evident that our customized neural network, as shown in Table 1, performs poorly. Several factors could contribute to the poor performance of the customized neural network. First, the high number of parameters significantly increases the risk of overfitting, particularly on a relatively small dataset like FashionMNIST. Second, the complex architecture, with multiple convolutional and fully connected layers, may lead to optimization challenges such as vanishing or exploding gradients, especially in deeper layers. Third, the lack of advanced regularization techniques, such as dropout or weight decay, might exacerbate the model’s inability to balance learning and generalization. Fourth, the computational demands of the large model might lead to slower convergence, requiring more training epochs and careful hyperparameter tuning. Based on the results, the primary reason for the poor performance appears to be vanishing gradients or the slower convergence. When the MLP classifier is replaced with the GLL classifier while keeping the same network structure, the performance improves significantly with the same feature encoder architecture. More details are discussed below.
We summarize the main takeaways of this experiment:
@@
    Accuracy comparison: In Table 3, networks with our GLL classifier have better accuracy than the MLP classifier under the same feature encoder structure, where the MLP classifier can be training directly or under the WNLL strategy. On the CIFAR-10 dataset, using ResNet architectures, replacing the MLP with GLL yields a modest improvement in accuracy. On the FashionMNIST dataset, where we employ our custom network, the improvement is more substantial. Our MLP accuracy on the CIFAR-10 dataset is similar to the WNLL accuracy, which is higher than the vanilla training results reported in the WNLL paper (Wang and Osher (2021)). One possible reason is that we incorporated SimCLR pretraining during the network training process, which provided a better initialization for the subsequent fully supervised training, resulting in improved performance.
@@
    Difference between MLP and GLL loss and accuracy in the MLP training: According to the FashionMNIST training curves (Figure 4), for MLP training (blue and orange curves), we provide both the MLP and GLL loss and accuracy during the training process. We observe that under the same feature encoder, the output of GLL corresponds to lower loss values and higher accuracy. This indicates that even if the information from GLL does not participate in the training, directly replacing MLP with GLL can achieve a significant improvement in performance. One reason is that the MLP classifier contains a large number of trainable parameters, which do not have a good classification effect when not well-trained. Another reason is that, even when well-trained, an MLP classifier consisting of several FC layers and ReLU activation functions can be understood as a series of hyperplanes that partition the feature space into classification regions (Montufar et al. (2014); Goodfellow et al. (2016)). While this approach can be effective, it may not fully capture the intrinsic structure of the data. In contrast, GLL’s Laplace Learning leverages the manifold structure of the data to guide the classification process. By considering the relationships between data points and their neighbors, GLL can potentially learn more complex and non-linear decision boundaries that better align with the underlying data distribution. This allows GLL to achieve superior classification performance compared to the hyperplane-based approach of MLP classifiers.

### v2

@@ 3.2.1 Sparse and efficient gradient computations
4 Implementation & Numerical Experiments
This section details the practical use of the proposed GLL classifier and showcases its strong performance through a variety of numerical experiments. Section 4.1 explains our implementation of the GLL as a classifier on supervised and semi-supervised tasks, including batching, contrastive pretraining, and data augmentation. Section 4.2 visualizes the embeddings learned by a GLL on a toy example and illustrates the effect of the Tikhonov parameter $\tau$ in (15); all remaining experiments use the standard Laplace learning algorithm in (11). Section 4.3 considers a deliberately over-parameterized CNN on FashionMNIST to compare a GLL head to an MLP head in terms of generalization, training dynamics, and computational efficiency. Section 4.4 presents large-scale experiments on the CIFAR-10 (krizhevsky2009learning) and EMNIST (cohen2017emnist) datasets at varying levels of supervision, comparing MLP, WNLL (wang2021graph), and GLL classification heads across several deep neural network backbones; GLL-based networks consistently match or surpass the performance of MLP- and WNLL-based networks, especially at low label rates. Section 4.5 quantifies the computational overhead of GLL by reporting training and inference times as a function of batch size and network size. Finally, Section 4.6 provides our most significant empirical result: across datasets, training strategies, and attack strengths, a GLL yields consistently and substantially stronger robustness to adversarial attacks than both MLP- and WNLL-based classifiers. While our theoretical analysis applies to general graph Laplacian-based learning algorithms, we leave numerical experiments on variants such as Poisson learning and $p$-Laplace learning to future work.
4.1 Implementation
In this subsection, we present implementation details for the GLL in our experiments. Although our GLL can be plugged into any neural network as a classifier, replacing the original head in this way leads to a batch setting that differs slightly from the classical MLP-based training paradigm. In addition, to accelerate training, we adopt a contrastive-learning–based pretraining stage, after which we perform supervised training for the neural network with the GLL.
4.1.1 Sampling and Training on Batches
We describe a slight modification of batching needed for a GLL classifier. Recalling the notation of Section 2.3, we let $\mathcal{X}$} denote the set of samples in the training set, $\mathcal{L}$ be the subset of base samples with known labels, $\mathcal{U}$ the unlabeled subset. Then we consider training on a batch $\mathcal{B}$ with size $B$ for the classic classifier (e.g. an MLP) or for our GLL classifier:
Classic Training:
Since each data is classified individually, the batch is selected only from the labeled dataset $\mathcal{L}$, i.e. $\mathcal{B}\subset\mathcal{L}$. The training loss $\mathcal{J}$ is computed between the predicted labels and their ground-truth labels over all samples in $\mathcal{B}$, which are available for every element of $\mathcal{B}$ since $\mathcal{B}\subset\mathcal{L}$.
GLL Training:
The data is classified by label propagation on a graph built on a batch $\mathcal{B}$. Because the GLL is a transductive classifier, it requires constructing each batch to include an additional set of labeled data - which we call the base dataset $\mathcal{L}_{b}$ - to propagate the labels to the rest of the batch. $\mathcal{L}_{b}$ is a subset sampled from the labeled dataset $\mathcal{L}$ with a size $N_{b}$. The remainder of the batch includes standard training samples $\mathcal{B}_{l}$ - whose labels we predict and compute a loss over - and (optionally) some unlabeled samples $\mathcal{B}_{u}$ in (e.g.) the semi-supervised setting. The batch can thus be written as
$\mathcal{B}=\mathcal{L}_{b}\cup\mathcal{B}_{l}\cup\mathcal{B}_{u},$ (50)
where $\mathcal{B}_{l}=(\mathcal{B}\setminus\mathcal{L}_{b})\cap\mathcal{L}$ and $\mathcal{B}_{u}=(\mathcal{B}\setminus\mathcal{L}_{b})\cap\mathcal{U}$. Let $N_{l}=|\mathcal{B}_{l}|$, $N_{u}=|\mathcal{B}_{u}|$, and $N_{b}+N_{u}=B$. We further require that, after fixing the base dataset $\mathcal{L}_{b}$, the proportion of labeled and unlabeled samples in each batch reflects their proportion in the remaining training set. Excluding the base samples, we sample $\mathcal{B}_{l}$ from $\mathcal{L}\setminus\mathcal{L}_{b}$ and $\mathcal{B}_{u}$ from $\mathcal{U}$ such that
$\frac{N_{l}}{N_{l}+N_{u}}=\frac{\left|\mathcal{L}\setminus\mathcal{L}_{b}\right|}{\left|\mathcal{L}\setminus\mathcal{L}_{b}\right|+\left|\mathcal{U}\right|}.$ (51)
Equivalently, for a fixed batch size $B$ and the base dataset $\mathcal{L}_{b}$, we choose $N_{l}$ and $N_{u}$ to satisfy $N_{l}+N_{u}=B$ and
$N_{l}=\frac{\left|\mathcal{L}\setminus\mathcal{L}_{b}\right|}{\left|\mathcal{L}\setminus\mathcal{L}_{b}\right|+\left|\mathcal{U}\right|}B,\quad N_{u}=\frac{\left|\mathcal{U}\right|}{\left|\mathcal{L}\setminus\mathcal{L}_{b}\right|+\left|\mathcal{U}\right|}B.$ (52)
In practice, we choose a suitable batch size $B$ and base dataset size $N_{b}$ such that the resulting $N_{l}$ and $N_{u}$ computed from (52) are both integers. For GLL, the training loss $\mathcal{J}$ is computed between the predicted labels and ground-truth labels over all samples in $\mathcal{B}_{l}$. Note that although the base dataset $\mathcal{L}_{b}$ is labeled, those labels are already used for graph-based label propagation and are therefore not included in the loss computation.
Remark 11 (Practical usage of GLL)
The proposed GLL-based batching scheme leads to several useful regimes in practice:
  1. 1.
    Semi-supervised learning. In the standard semi-supervised setting, the training set decomposes as ${\mathcal{X}}=\mathcal{L}\cup\mathcal{U}$ with $\mathcal{L}$ labeled and $\mathcal{U}$ unlabeled. In each batch, the nodes in $\mathcal{B}_{u}$ are treated as unlabeled on the graph: they do not contribute directly to the loss, but they participate in graph construction and label propagation, and hence influence the predictions on $\mathcal{B}_{l}$. In this way the GLL can fully exploit the unlabeled data to improve classification on the labeled subset.
  2. 2.
    Choice of base dataset during training. During training, the base dataset $\mathcal{L}_{b}$ is selected at the beginning of each epoch, and all batches within the same epoch share the same base set $\mathcal{L}_{b}$. We consider two strategies for constructing $\mathcal{L}_{b}$, including a simple stratified random resampling scheme that preserves the class proportions of the full labeled set $\mathcal{L}$ and an uncertainty-based selection method. The precise sampling procedures are described in Appendix A.1.
  3. 3.
    Transductive inference at test time. At test time, the GLL still requires a base set to construct the graph, but no backpropagation through the GLL classifier is needed. We can therefore first apply the feature encoder (the network backbone preceding the GLL classifier) to all available samples, including both training and test data, using standard mini-batch processing. Based on the resulting feature representations, we then construct a single large graph and solve the graph learning problem described in Section 2.3 on this graph to obtain predictions for the test nodes and compute the test accuracy. Since this stage only involves linear algebra on fixed features and does not require gradients, it can be moved to the CPU to reduce GPU memory usage.
4.1.2 Pretraining
Due to its transductivity, GLL accuracy and learning behavior may degenerate if the feature embeddings are poor, which is often the case for an untrained network. To remedy this, we warm the model up with a pretext task that allows the feature extractor to learn basic features. Prior works (brown2023utilizing; chen2025cgap) show that feature representations learned from contrastive learning pre-training exhibit more pronounced clustering structure, leading to improved performance for graph-based learning methods such as GLL. We follow suit and warm-up the models using the contrastive methods SimCLR (chen2020simple) or SupCon (khosla2020supervised), depending on the availability of labeled data. Contrasive learning and the GLL are a natural pairing, as the loss functions for the contrastive methods use a similarity function that is effectively the same as the one used in the similarity graph construction in the GLL. Warming up the models with contrastive learning will thus help the model capture features and structures similar to the GLL and accelerate learning. The contrastive methods are discussed further in Appendix A.2. Warm-ups were also used in previous work on graph-based classification heads (wang2021graph), where the authors alternate between 400 epochs of MLP-based training and 5 epochs of WNLL-based training in order to stabilize learning. Conversely, we warm-up with contrastive learning, and then proceed with GLL training, with no alternating stages.
In practice, we consider a mixed training target with the SimCLR loss $\mathcal{L}_{\mathrm{sim}}$ (59) and the SupCon loss $\mathcal{L}_{\mathrm{sup}}$ (60). We only pretrain the feature encoder before the classifier in the neural network architecture since both losses do not require predicted labels. The pretraining loss is defined by:
$\mathcal{L}_{\mathrm{pretrain}}=\gamma\mathcal{L}_{\mathrm{sim}}+(1-\gamma)\mathcal{L}_{\mathrm{sup}},\gamma\in[0,1].$ (53)
The contrastive pretraining described above is agnostic to the choice of classification head and can therefore be used both with classical MLP classifiers and with our GLL classifier. For each pretraining batch, we draw a mixture of labeled and unlabeled samples from the full training set $\mathcal{X}=\mathcal{L}\cup\mathcal{U}$, in the same spirit as the batch construction in Section 4.1.1. The SimCLR loss $\mathcal{L}_{\mathrm{sim}}$ is evaluated on all samples in the batch, regardless of whether they are labeled or unlabeled, since it does not require label information. By contrast, the SupCon loss $\mathcal{L}_{\mathrm{sup}}$ is computed only on those samples in the pretraining batch that have labels, using their ground-truth class assignments. In this way, the pretraining stage can exploit all available data while still benefiting from supervised signal when labels are present, and the resulting encoder can be paired with any downstream classifier, including GLL.
Remark 12 (Choice of mixing parameter $\gamma$)
In practice, we choose the mixing parameter $\gamma$ in $\mathcal{L}_{\mathrm{pretrain}}$ (53) from the discrete candidate set $\{0.01,0.25,0.5,0.75,0.99\}$. For each candidate value, we perform the contrastive pretraining stage to obtain a feature encoder and then feed the resulting feature vectors into a parameter-free graph learning classifier. We compute the resulting classification accuracy on the labeled subset $\mathcal{L}\subset\mathcal{X}$ of the training data and select the value of $\gamma$ that achieves the highest accuracy on this labeled subset.
4.1.3 Data Augmentation
Data augmentation is applied throughout both the contrastive pretraining stage without a classification head and the subsequent supervised training of the full network with different classifiers. This follows the standard practice in contrastive representation learning (chen2020simple; chen2020big; khosla2020supervised), where multiple views of the same input are generated via stochastic transformations to form positive pairs. During evaluation on all benchmarks, we disable data augmentation and feed the original inputs to the encoder and classifier.
Concretely, at each training step we first sample a mini-batch of images from the training set and then apply a stochastic augmentation pipeline to each image. During contrastive pretraining, every sample in the batch is independently transformed twice to produce two augmented views, which are then used as positive pairs for the SimCLR and SupCon losses. In the fully supervised training phase, each data sample is augmented once on-the-fly to increase the diversity of training examples.
The augmentation pipeline itself is based on a RandAugment-style (cubuk2019randaugment) policy: for each image, a small number of transformations is drawn at random from a pool that includes common photometric and geometric operations, such as autocontrast, histogram equalization, brightness/contrast adjustments, rotations, shears, translations, and solarization, followed by a random cutout operation that masks out a square region of the image. For grayscale datasets we use a slightly modified variant of this policy that is tailored to single-channel images, e.g., removing color-only operations and using grayscale cutout. The full list of transformations and the exact sampling procedure are provided in Appendix A.3. RandAugment is applied to all samples in the batch, including those belonging to the base dataset $\mathcal{L}_{b}$.
@@ 4.1.3 Data Augmentation
Figure 3: The initial embedding of the two moons dataset from the random seed. The blue and red correspond to the different classes, with the starred samples corresponding to the base samples for the respective classes.
4.2 Understanding Embeddings and the Effect of Tikhonov Regularization via Two Moons
In this subsection, we highlight the differences between the learned embeddings of neural networks when trained with a GLL compared to a softmax layer. Moreover, we demonstrate the effect on the embedding by modifying the decay parameter, $\tau$, from Section 2.4.
The softmax classifier uses no relational information and thus will create an embedding that leads to a straightforward separation based on the features alone. The GLL, using relational information, will learn to separate the data such that the generated similarity graph will be nearly disjoint by class (a totally disjoint graph can prevent learning). Therefore, we may expect the GLL to produce a more relaxed embedding, as a disjoint similarity graph does not necessarily require the data to be separated to a probability simplex. Increasing the Tikhonov regularization, $\tau$, inhibits label propagation and weakens predictions, meaning higher confidence predictions and lower loss will require a similarity graph that is even more separated. Specifically, increasing $\tau$ will increase the learning rate and convergence speed of the process and should lead to more locally dense learned embeddings. We give a brief example of these effects in two dimensions with the simple toy dataset, two moons (scikit-learn)).
In this experiment, we perform supervised classification on a two-dimensional generated dataset of 500 samples from the two moons distribution with a noise of 0.1 and random state of one. The network was trained with a stochastic gradient descent optimizer with a learning rate of 0.01.
The architecture backbone for this experiment is a simple multi-layer perceptron neural network with three fully connected (FC) layers and two ReLU activation layers. The hidden layer has a size of 64 and the output size is two for two dimensional visualization purposes. We choose an initialization for the neural network such that the initial embedding has sufficiently mixed samples (Figure 3). As the dataset is simplistic, there will be no projection head in this comparison, meaning the only difference between the standard MLP model and the GLL model will be purely the softmax and GLL classifiers.
We arbitrarily fix the first ten samples, five from each class, to act as the base samples which propagate the labels. As the dataset is small, we load all the samples in each batch so there is no concern about partial class representation. To make the results more interpretable, we fix the $\varepsilon_{k}$ term, from the similarity equation, Equation 1, to be fixed at one.
The embeddings for this ablation study are shown in Figure 4. We compare five values of $\tau$: 0, 0.001, 0.01, 0.1, and 0.5. The figure shows the learned embedding of the dataset from the neural network across lengthening training epochs.
We note that this regime is deliberately constrained to two dimensions, meaning that the network may struggle to ’move’ the embeddings from one class beyond the other class’s base samples, as approaching the opposing base samples will increase the loss. The seed for these samples was chosen to be difficult and we note that the problem is easier in a higher dimensional space. Despite this, after sufficient training, the model successfully separates the data.
@@ 4.2 Understanding Embeddings and the Effect of Tikhonov Regularization via Two Moons
Figure 4: Visualization of effect of $\tau$ on spatial embeddings. Each row corresponds to a graph learning model with a specified value of $\tau$ or a softmax classification head. The columns correspond to the epoch of the training. With low values of $\tau$, the model separates the data by class and resembles the original two moons shape. Increasing $\tau$ forces tighter clustering and localization. The softmax layer simply learns a linear separation with no regard to relational information. Further training with the softmax layer increases the scale, which is not seen without the axes.
Examining the results, the baseline MLP encoder trained with the softmax layer simply separated the data approximately along a line. However, the same encoder architecture trained with the GLL will recover the original two moons shape, implying that the two moons shape induces a strongly separated $k$-nearest neighbors similarity graph.
The addition of the $\tau$ decay parameter will encourage a tighter clustering of the classes around the base labels. For low values of $\tau$, we still see initial convergence to the two moons before diverging into tighter clusters around the base samples. If the $\tau$ value is large, the divergence is more immediate and is seen splintering one of the classes into multiple clusters localized about different base samples. This can be rationalized as the base samples pulling the unlabeled samples closer to themselves, strengthening their own influence and dampening the influence and learning from distant base samples. This is because increasing $\tau$ acts as an exponential dampening term which localizes the influence of the base samples.
4.3 Toy Example: Over-parameterized CNN on FashionMNIST
In this subsection, we focus on a toy example on the FashionMNIST dataset to compare our GLL classifier with a MLP classifier. The goal is to study a setting where the backbone network is intentionally over-parameterized and challenging to optimize, so that the effect of replacing the MLP classifier by GLL becomes more evident.
Experimental setup.
We use the FashionMNIST dataset (xiao2017fashion) and construct a simple convolutional neural network as the feature encoder. The architecture is summarized in Table 1. The feature encoder consists of three convolutional layers followed by max-pooling layers and a fully connected (FC) layer. Each convolutional and FC layer is followed by a ReLU activation function, which is omitted from the table for brevity. The resulting encoder is purposely over-parameterized relative to FashionMNIST, and we will see that it is difficult to train effectively when combined with a standard MLP classifier.
@@ 4.3 Toy Example: Over-parameterized CNN on FashionMNIST
Table 1: Network structure used in the FashionMNIST experiments (Section 4.3). The feature encoder is deliberately over-parameterized to create a challenging training scenario, and the model contains 1,975,424 parameters in the feature encoder and 142,346 parameters in the MLP classifier.
The MLP classifier consists of two fully connected layers followed by a softmax layer, as shown in Table 1. To compare with our GLL classifier, we keep the feature encoder architecture fixed and replace the MLP classifier by GLL, which has no trainable parameters. In all cases, the final outputs are used to compute the cross-entropy loss with the ground-truth labels. All experiments in this subsection are fully supervised and use the entire FashionMNIST training set.
Training strategies.
We consider four training strategies on FashionMNIST, all sharing the same feature encoder:
  • MLP only (100 MLP). We train the network with the MLP classifier from scratch for 100 epochs. During this process, the encoder output is also passed to a GLL classifier in parallel. Since GLL has no trainable parameters, this does not affect the training dynamics. Throughout MLP training, we use the GLL outputs to compute loss (without backpropagation) and test error, allowing a direct comparison between MLP and GLL under the same encoder.
  • GLL-0 (50 GLL). We train the network from scratch using GLL as the classifier for 50 epochs. In the plots we refer to this strategy as GLL-0.
@@
  • GLL-50 (50 MLP + 50 GLL). We first train the network with the MLP classifier for 50 epochs. Starting from epoch 51, we switch the classifier to GLL and continue training for another 50 epochs, using the gradients induced by the GLL loss.
  • GLL-75 (75 MLP + 25 GLL). We first train with the MLP classifier for 75 epochs, and then switch to GLL for the remaining 25 epochs.
Test error comparison.
Table 2 reports the final test classification errors in %. For the MLP-only strategy, we report the test error of the MLP classifier, which is used for training, and the test error of the GLL classifier evaluated on the same encoder. For the remaining strategies, we report the test error of the GLL classifier at the end of training.
Dataset Trial Epochs Test error (%)
FashionMNIST MLP only (MLP) 100 MLP 45.42
FashionMNIST MLP only (GLL) 100 MLP 14.66
FashionMNIST GLL only (GLL-0) 50 GLL 8.90
FashionMNIST GLL-50 50 MLP + 50 GLL 8.68
FashionMNIST GLL-75 75 MLP + 25 GLL 8.91
Table 2: Test classification error rates (%, lower is better) on FashionMNIST for the customized network in Table 1 under different training strategies. For the “MLP only” strategy we report both the error of the MLP classifier and that of a GLL classifier evaluated on the same encoder (GLL). For the remaining strategies, GLL is used as the classifier during training and evaluation.
Figure 5 shows the evolution of the training loss and test accuracy over epochs. Figure 5(a) presents the training loss curves for all four strategies, while Figure 5(b) presents the corresponding test accuracies. The curves for GLL-50 and GLL-75 start at epochs 50 and 75, respectively, since these strategies switch to GLL after an initial MLP training phase.
@@
Figure 5: Training loss and test accuracy of four training strategies on the FashionMNIST dataset with the customized feature encoder in Table 1. The strategies are: (1) MLP trained from scratch, (2) GLL trained from scratch (GLL-0), (3) GLL-50 (GLL after 50 epochs of MLP training), and (4) GLL-75 (GLL after 75 epochs of MLP training). For the MLP-only strategy, both the MLP and GLL losses and accuracies are recorded during training; for the remaining strategies, only the GLL classifier is used.
From Table 2, the MLP classifier trained from scratch performs poorly: after 100 epochs it still incurs a test error of 45.42%, even though the GLL classifier reaches a much lower test error of 14.66% when evaluated on the same encoder. This suggests that the bottleneck is not the representation learned by the encoder, but rather the optimization of the over-parameterized MLP classifier. In contrast, when we replace the MLP with the parameter-free GLL classifier and train the same backbone (GLL-0), the test error decreases to 8.90%.
The mixed strategies GLL-50 and GLL-75 behave similarly. Although the MLP training tends to stagnate, switching to GLL mid-training yields stable decreases in the loss and further improvements in test error, reaching 8.68% and 8.91%, respectively. This indicates that once the encoder has learned a reasonable representation, GLL can immediately exploit it more effectively than the MLP classifier, and can also continue to drive smooth optimization of the encoder.
We summarize the main observations from this toy example:
@@
    Error under a fixed encoder. Even when the encoder is trained only with the MLP classifier, the GLL achieves substantially lower loss and lower test error than the MLP on the same features. Simply replacing the MLP classifier by GLL can therefore yield a large performance gain without modifying the encoder architecture.
@@
    Robustness to over-parameterization of the classifier. The over-parameterized MLP classifier is difficult to optimize and can suffer from issues such as vanishing gradients or dying ReLUs, leading to training stagnation. The GLL classifier, having no trainable parameters, avoids these optimization pathologies and works well even when the backbone is not carefully designed.